# Preserving Historical Truth: Detecting Historical Revisionism in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are increasingly used as sources of historical information, motivating the need for scalable audits on contested events and politically charged narratives in settings that mirror real user interactions. We introduce HistoricalMisinfo, a curated dataset of $500$ contested events from $45$ countries, each paired with a factual reference narrative and a documented revisionist reference narrative. To approximate real-world usage, we instantiate each event in $11$ prompt scenarios that reflect common communication settings (e.g., questions, textbooks, social posts, policy briefs). Using an LLM-as-a-judge protocol that compares model outputs to the two references, we evaluate LLMs varying across model architectures in two conditions: (i) neutral user prompts that ask for factually accurate information, and (ii) robustness prompts in which the user explicitly requests the revisionist version of the event. Under neutral prompts, models are generally closer to factual references, though the resulting scores should be interpreted as reference-alignment signals rather than definitive evidence of human-interpretable revisionism. Robustness prompting yields a strong and consistent effect: when the user requests the revisionist narrative, all evaluated models show sharply higher revisionism scores, indicating limited resistance or self-correction. HistoricalMisinfo provides a practical foundation for benchmarking robustness to revisionist framing and for guiding future work on more precise automatic evaluation of contested historical claims to ensure a sustainable integration of AI systems within society.
11footnotetext: Code and Dataset:  francescortu/PreservingHistoricalTruth
@@ 1 Introduction
LLMs present new challenges in this context. Trained on vast and often uncurated corpora, they risk reproducing inaccuracies, biases, and deliberate falsehoods at scale (Bender et al., 2021; Bommasani et al., 2021; Stammbach et al., 2024). For instance, recent studies have shown that LLMs often display WEIRD (Western, Educated, Industrialized, Rich, Democratic) biases Mihalcea et al. (2025), stemming from their training data, model architectures, and evaluation frameworks. These biases inherently privilege the historical narratives of dominant groups while marginalizing alternative perspectives (Santurkar et al., 2023; Ryan et al., 2024). Beyond inherited bias, highly centralized control over training data and fine-tuning introduces further risks: the small set of organizations developing LLMs can, in principle, shape outputs to reflect preferred narratives or suppress inconvenient facts. As these systems increasingly mediate public access to history, understanding and auditing their susceptibility to revisionism is critical.
@@
  • RQ1 Under neutral prompts, how do model responses align with factual versus revisionist reference narratives across countries and context?
  • RQ2 How does this alignment signal vary across different user interaction scenarios?
  • RQ3 How robust are LLMs when directly prompted to generate revisionist content?
@@ 2 Related Work
Historical revisionism presents a significant threat to social justice and information integrity, especially under authoritarian regimes that manipulate history to reinforce propaganda. Social science research has identified systematic mechanisms, such as political repression, educational revisionism, social engineering, cultural erasure, or media censorship, as tools for shaping regime-aligned historical narratives Belmonte and Rochlitz (2020); Hahn and Hahn (2005); Boyce and O’Day (1996); Kopeček (2008); Kasianov (2011). Causal frameworks describe how regimes reconstruct national histories to secure political legitimacy, often by invoking collective memories of past turmoil Belmonte and Rochlitz (2020); Kasianov (2011); Boyce and O’Day (1996). These reconstructions align historical interpretation with contemporary ideological goals, particularly in post-Soviet and post-colonial contexts Boyce and O’Day (1996); Kasianov (2011). Based on this, studies in information systems show that digital platforms, especially social networks, have become channels for revisionist narratives. For example, coordinated campaigns and bot networks are frequently used to manipulate historical discourse during geopolitical crises such as the Russian invasion of Ukraine Geissler et al. (2023). As LLMs become prominent tools for information retrieval Kasneci et al. (2023), their influence over public understanding of history raises new concerns. LLMs may inherit biases from politically influenced training data or inadvertently replicate revisionist narratives.
@@ 4 LLM Testing Pipeline
To systematically evaluate how LLMs respond to historically contested content, we design a three-stage testing pipeline. First, we construct prompt scenarios that reflect the channels through which revisionist narratives are typically propagated (4.1). Second, we run LLM inference across multiple medium-sized models (4.2). Third, we assess model outputs using an LLM-as-a-judge (Zheng et al., 2023) framework to measure the presence of revisionist or factual content (4.3). This design moves beyond single-prompt evaluation by grounding assessment in real-world usage scenarios and a reference-based comparison framework (factual vs. revisionist narratives), enabling systematic and reproducible measurement of how model outputs align with competing accounts rather than assigning absolute revisionism labels.
@@ 5 Results
This section presents the main findings of our evaluation. We begin with descriptive analyses, highlighting how LLMs respond across prompt scenarios, historical topics, and geographic regions.
@@ Acknowledgment
We thank Jonathon Penney for helpful discussions and suggestions. This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by Schmidt Sciences SAFE-AI Grant; by NSERC Discovery Grant RGPIN-2025-06491; MPI funding: https://atlas.is.localnet/confluence/display/SCO/Affiliation+and+Acknowledgements+in+Publications#AffiliationandAcknowledgementsinPublications-ClusterofExcellenceMachineLearning:NewPerspectivesforScience,UniversityofT%C3%BCbingen by a National Science Foundation award (#2306372); by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung. The usage of OpenAI credits is largely supported by the Tübingen AI Center. Francesco Ortu and Alberto Cazzaniga are supported by the Italian region of Friuli-Venezia Giulia (CUP:F53C22001770002).

### v2

@@ Abstract
Large language models (LLMs) are increasingly used as sources of historical information, motivating the need for scalable audits on contested events and politically charged narratives in settings that mirror real user interactions. We introduce HistoricalMisinfo, a curated dataset of $500$ contested events from $45$ countries, each paired with a factual reference narrative and a documented revisionist reference narrative. To approximate real-world usage, we instantiate each event in $11$ prompt scenarios that reflect common communication settings (e.g., questions, textbooks, social posts, policy briefs). Using an LLM-as-a-judge protocol that compares model outputs to the two references, we evaluate LLMs varying across model architectures in two conditions: (i) neutral user prompts that ask for factually accurate information, and (ii) robustness prompts in which the user explicitly requests the revisionist version of the event. Under neutral prompts, models are generally closer to factual references, though the resulting scores should be interpreted as reference-alignment signals rather than definitive evidence of human-interpretable revisionism. Robustness prompting yields a strong and consistent effect: when the user requests the revisionist narrative, all evaluated models show sharply higher revisionism scores, indicating limited resistance or self-correction. HistoricalMisinfo provides a practical foundation for benchmarking robustness to revisionist framing and for guiding future work on more precise automatic evaluation of contested historical claims to ensure a sustainable integration of AI systems within society. [^1] Code and dataset:  francescortu/PreservingHistoricalTruth
@@ 1 Introduction
LLMs present new challenges in this context. Trained on vast and often uncurated corpora, they risk reproducing inaccuracies, biases, and deliberate falsehoods at scale (Bender et al., 2021; Bommasani et al., 2021; Stammbach et al., 2024). For instance, recent studies have shown that LLMs often display WEIRD (Western, Educated, Industrialized, Rich, Democratic) biases (Mihalcea et al., 2025), stemming from their training data, model architectures, and evaluation frameworks. These biases inherently privilege the historical narratives of dominant groups while marginalizing alternative perspectives (Santurkar et al., 2023; Ryan et al., 2024). Beyond inherited bias, highly centralized control over training data and fine-tuning introduces further risks: the small set of organizations developing LLMs can, in principle, shape outputs to reflect preferred narratives or suppress inconvenient facts. As these systems increasingly mediate public access to history, understanding and auditing their susceptibility to revisionism is critical.
@@
  • RQ1 Under neutral prompts, how do model responses align with factual versus revisionist reference narratives across countries and context? (Sec. 5.1)
  • RQ2 How does this alignment signal vary across different user interaction scenarios? (Sec.5.2)
  • RQ3 How robust are LLMs when directly prompted to generate revisionist content? (Sec.5.3)
@@ 2 Related Work
Historical revisionism presents a significant threat to social justice and information integrity, especially under authoritarian regimes that manipulate history to reinforce propaganda. Social science research has identified systematic mechanisms, such as political repression, educational revisionism, social engineering, cultural erasure, or media censorship, as tools for shaping regime-aligned historical narratives (Belmonte and Rochlitz, 2020; Hahn and Hahn, 2005; Boyce and O’Day, 1996; Kopeček, 2008; Kasianov, 2011). Causal frameworks describe how regimes reconstruct national histories to secure political legitimacy, often by invoking collective memories of past turmoil (Belmonte and Rochlitz, 2020; Kasianov, 2011; Boyce and O’Day, 1996). These reconstructions align historical interpretation with contemporary ideological goals, particularly in post-Soviet and post-colonial contexts (Boyce and O’Day, 1996; Kasianov, 2011). Based on this, studies in information systems show that digital platforms, especially social networks, have become channels for revisionist narratives. For example, coordinated campaigns and bot networks are frequently used to manipulate historical discourse during geopolitical crises such as the Russian invasion of Ukraine (Geissler et al., 2023). As LLMs become prominent tools for information retrieval (Kasneci et al., 2023), their influence over public understanding of history raises new concerns. LLMs may inherit biases from politically influenced training data or inadvertently replicate revisionist narratives.
@@ 4 LLM Testing Pipeline
To systematically evaluate how LLMs respond to historically contested content, we design a three-stage testing pipeline. First, we construct prompt scenarios that reflect the channels through which revisionist narratives are typically propagated (Sec. 4.1). Second, we run LLM inference across multiple medium-sized models (Sec. 4.2). Third, we assess model outputs using an LLM-as-a-judge (Zheng et al., 2023) framework to measure the presence of revisionist or factual content (Sec. 4.3). This design moves beyond single-prompt evaluation by grounding assessment in real-world usage scenarios and a reference-based comparison framework (factual vs. revisionist narratives), enabling systematic and reproducible measurement of how model outputs align with competing accounts rather than assigning absolute revisionism labels.
@@ 5 Results
This section presents the main findings of our evaluation. We begin with descriptive analyses, and conduct further investigations of how LLMs respond across prompt scenarios, historical topics, and geographic regions.
@@ Acknowledgment
We thank Jonathon Penney for helpful discussions and suggestions. Francesco Ortu and Alberto Cazzaniga are supported by the Italian region of Friuli-Venezia Giulia (CUP:F53C22001770002). This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by the Survival and Flourishing Fund; by Coefficient Giving; and by the Cooperative AI Foundation. The usage of OpenAI credits is largely supported by the Tübingen AI Center and Schmidt Sciences. Resources used in preparing this research project were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.